solve the system by addition
2x - 9y = -4
x - 4y = -1

Answer by scott8148(6628)

multiply the second equation by -2 to get -2x+8y=2 ... add this result to the first equation to get

(2x-9y)+(-2x+8y)=-4+2 ... -y=-2 ... y=2 ... x-4(2)=-1 ... x=7